# DR-RAG: Applying Dynamic Document Relevance to Retrieval-Augmented Generation for Question-Answering

### v1

@@ A.4 Baseline
A.5 equation
A.5.1 QDC
$\begin{split}Context&=\{\}\\ \{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{k_{1}}\}&=\texttt{Retriever}(\boldsymbol{q})\\ Context&=Context\cup\{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{k_{1}}\}\\ \boldsymbol{q_{m}^{*}}&=\texttt{Concat}(\boldsymbol{q},\boldsymbol{d}_{i})\\ \{\boldsymbol{d}^{\prime}_{1},\boldsymbol{d}^{\prime}_{2},\ldots,\boldsymbol{d}^{\prime}_{k}\}&=\texttt{Retriever}(\boldsymbol{q_{m}^{*}})\\ Context&=Context\cup\{\boldsymbol{d}^{\prime}_{j}\mid\boldsymbol{d}^{\prime}_{j}\not\in Context\text{ and }first\}\\ answer&=\texttt{LLM}(\texttt{Concat}(\boldsymbol{q},Context))\end{split}$ (3)
A.5.2 CIS
$\text{Context}=\{\}$
$\{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{k_{1}}\}=\texttt{Retriever}(\boldsymbol{q})$
$\text{Context}=\text{Context}\cup\{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{\frac{k}{2}}\}$
$\text{new}_{\text{query}}(i)=\texttt{Concat}(\boldsymbol{q},\boldsymbol{d}_{i}),\quad i=1,2,\ldots,\frac{k}{2}$
$\{\boldsymbol{d}^{\prime}_{1},\boldsymbol{d}^{\prime}_{2},\ldots,\boldsymbol{d}^{\prime}_{k}\}=\texttt{Retriever}(\text{new}_{\text{query}}(i)),\quad i=1,2,\ldots,\frac{k}{2}$
$\text{Context}=\text{Context}\cup\{\boldsymbol{d}^{\prime}_{j}\mid\boldsymbol{d}^{\prime}_{j}\not\in\text{Context}\text{ and }\text{first}\},\quad i=1,\ldots,\frac{k}{2}$
$=\{\boldsymbol{d}_{i}\mid\exists j\neq i,C(\boldsymbol{d}_{i},\boldsymbol{d}_{j})\neq 0,\quad\boldsymbol{d}_{i},\boldsymbol{d}_{j}\in\text{Context}\}$ Context }
@@ A.5.2 CIS
A.5.3 CFS
$\text{Context}=\{\}$
$\{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{\frac{k}{2}}\}=\texttt{Retriever}(\boldsymbol{q})$
$\text{Context}=\text{Context}\cup\{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{\frac{k}{2}}\}$
$\text{new}_{\text{query}}(i)=\texttt{Concat}(\boldsymbol{q},\boldsymbol{d}_{i}),\quad i=1,2,\ldots,\frac{k}{2}$
$\{\boldsymbol{d}^{\prime}_{1},\boldsymbol{d}^{\prime}_{2},\ldots,\boldsymbol{d}^{\prime}_{k}\}=\texttt{Retriever}(\text{new}_{\text{query}}(i)),\quad i=1,2,\ldots,\frac{k}{2}$
$\text{Context}=\text{Context}\cup\{\boldsymbol{d}^{\prime}_{j}\mid\boldsymbol{d}^{\prime}_{j}\not\in\text{Context}\text{ and }C(\boldsymbol{q},\boldsymbol{d}_{i},\boldsymbol{d}^{\prime}_{j})=1\text{ and }\text{first}\},\quad i=1,\ldots,\frac{k}{2}$
@@ A.5.3 CFS
Table 7: Case of the classifier we train for select in CIS and CFS.
@@
Table 8: Cases the Query can be answered correctly in DR-RAG, and can not in Adaptive-RAG, where we present wrong answer in red and the right answer in blue.
@@
Table 9: A case of our prompt provided to LLMs.
documents
Document 1:
Games
Document 2:
with film melodramas like" Imitation of Life All That Heaven Allows Written on the WindMagnificent Obsession" and" A Time
to Love and a Time to Die". While those films were initially panned by critics as sentimental women’s pictures, they are today
widely regarded by film directors, critics and scholars as masterpieces. His work is seen as" critique of the bourgeoisie in general
and of 1950s America in particular", while painting a" compassionate portrait of characters trapped by social conditions". Beyond
the surface of the film, Sirk worked with complex mise en scenes and lush Technicolor colors to subtly underline his message.
Document 3:
Anne Curtis. The film is co-produced by Star Cinema and Viva Films under the working title" Momalland". The film premiered
in Philippine cinemas on December 25, 2019 as one of the official entries to the 2019 Metro Manila Film Festival. " The Mall,
The Merrier" marks the first on- screen collaboration between Anne Curtis and Vice Ganda, both of whom are regular hosts in
the noontime variety show" It’s Showtime".
Document 4:
He was nominated for two Academy Awards for Best Writing, Original Story and Best Writing, Screenplay on the 1943 film"
The More the Merrier". He died in 1992 in New York City, shortly after his 80th birthday.
Document 5:
Don Ameche.
Document 6:
during World War II, especially in Washington, D.C. The picture stars Jean Arthur, Joel McCrea and Charles Coburn. The movie
was directed by George Stevens. The film was written by Richard Flournoy, Lewis R. Foster, Frank Ross, and Robert Russell,
from" Two’s a Crowd", an original story by Garson Kanin( uncredited). This film was remade in 1966 as" Walk, Do n’t Run",
with Cary Grant, Samantha Eggar and Jim Hutton.
After reading the documents above, answering the following question. Reasoning step by step. At last, you should output the final
result via the following format:
Answer: <your answer based on the documents>;
Please answer the question directly.
question
Table 10: A case of documents retrieved by QDC and CFS, where the documents we need are in blue, and the top-$k$ is 4.
@@
Table 11: A case of documents retrieved by QDC and CIS, where the documents we need are in blue, and the top-$k$ is 4.
@@
Who is the spouse of the
Green performer?
Document 1:
@@
Table 12: Results of 500 samples sampled on MuSiQue based on gpt-3.5-turbo and gpt-4-turbo.
top-$k$ LLMs EM F1 Acc
3 gpt-3.5-turbo 26.00 36.20 35.00
gpt-4-turbo 33.80 46.30 47.20
4 gpt-3.5-turbo 25.80 37.60 38.60
gpt-4-turbo 34.20 47.80 51.00
6 gpt-3.5-turbo 28.40 41.10 40.60
gpt-4-turbo 39.40 52.70 54.20

### v3

@@ 5 Conclusion
6 Limitations
7 Ethics Statement
DR-RAG substantiates its efficacy in real-world scenarios, which are characterized by diverse user queries. However, given the potential variability in user inputs, which may span a range from benign to offensive, it is imperative to consider scenarios where inputs might be detrimental. Such instances could facilitate the retrieval of objectionable content and lead to unsuitable responses by retrieval-augmented LLMs. Addressing this concern necessitates the development of robust methodologies to detect and mitigate offensive or inappropriate content in both user inputs and the documents retrieved within the RAG framework. This area represents a critical part for future research.
@@ A.3 Baseline
A.4 Related Works
A.4.1 RAG for Multi-hop QA
A.4.2 Retriever in RAG
A.5 More Analysis
More Cases    Table 9 shows the prompts we provide to LLMs. Contexts contain the documents (Document $i$) after retrieving and selecting. Moreover, we show the case of the classifier for selection in CIS and CFS methods in Table 10, the output cases compared to Adaptive-RAG in Table 11, the case of documents retrieved by QDC and CFS methods in Table 12, the case of documents retrieved by QDC and CIS methods in Table 13.
Algorithm 1 Classifier Forward Selection (CFS)
1:
2:Classifier $C$
3:Retrieval Function Retriever
4:Input query $\boldsymbol{q}$
5:Generated response $answer$
@@ A.5 More Analysis
7:Retrieve $k_{1}$ documents: $\{\boldsymbol{d}_{1},\boldsymbol{d}_{2},\ldots,\boldsymbol{d}_{k_{1}}\}=\texttt{Retriever}(\boldsymbol{q})$
@@
9:for $i=1$ to $k_{1}$ do
10:     Construct a new query: $\boldsymbol{{q_{i}}^{*}}=\texttt{concat}(\boldsymbol{q},\boldsymbol{d_{i}})$
11:end for
12:Retrieve full set of documents for each new query:
[…]
14:for $i=1$ to $k_{1}$ do
15:     for $j=1$ to $k_{2}$ do
16:         if $\boldsymbol{d}^{\prime}_{i,j}\not\in\text{$Cnt$}$ and $C\text{(}\boldsymbol{q},\boldsymbol{d}_{i},\boldsymbol{d}^{\prime}_{i,j})=positive$ then
[…]
18:         end if
19:     end for
20:end for
[…]
23:return $answer$
Table 9: A case of our prompt provided to LLMs.
Contexts
Document 1:
Games
Document 2:
with film melodramas like "Imitation of Life All That Heaven Allows Written on the WindMagnificent Obsession" and "A
Time to Love and a Time to Die". While those films were initially panned by critics as sentimental women’s pictures, they are
today widely regarded by film directors, critics and scholars as masterpieces. His work is seen as "critique of the bourgeoisie
in general and of 1950s America in particular", while painting a" compassionate portrait of characters trapped by social con-
ditions". Beyond the surface of the film, Sirk worked with complex mise enscenes and lush Technicolor colors to subtly un-
derline his message.
Document 3:
Anne Curtis. The film is co-produced by Star Cinema and Viva Films under the working title" Momalland". The film pre-
miered in Philippine cinemas on December 25, 2019 as one of the official entries to the 2019 Metro Manila Film Festival. "
The Mall, The Merrier" marks the first on- screen collaboration between Anne Curtis and Vice Ganda, both of whom are
regular hosts in the noontime variety show" It’s Showtime".
Document 4::
He was nominated for two Academy Awards for Best Writing, Original Story and Best Writing, Screenplay on the 1943 film
"The More the Merrier". He died in 1992 in New York City, shortly after his 80th birthday.
Document 5:
Don Ameche.
Document 6:
during World War II, especially in Washington, D.C. The picture stars Jean Arthur, Joel McCrea and Charles Coburn. The
movie was directed by George Stevens. The film was written by Richard Flournoy, Lewis R. Foster, Frank Ross, and Robert
Russell, from" Two’s a Crowd", an original story by Garson Kanin( uncredited). This film was remade in 1966 as" Walk,
Don’t Run", with Cary Grant, Samantha Eggar and Jim Hutton.
After reading the documents above, answering the following question. Reasoning step by step. At last, you should output the
final result via the following format:
Answer: <your answer based on the documents>;
Please answer the question directly.
Question
Table 10: Case of the classifier we train for selection in CIS and CFS methods. We mark relevant information that can influence classification results in blue.
@@
Table 11: Cases that the query can be answered correctly in DR-RAG, and can not in Adaptive-RAG. We present wrong answer in red and the right answer in blue.
@@
Table 12: A case of documents retrieved by QDC and CFS on the MuSiQue dataset, where the necessary documents are in blue, and the top-$k$ is 4.
@@
Table 13: A case of documents retrieved by QDC and CIS on the HotpotQA dataset, where the necessary documents are in blue, and the top-$k$ is 4.
@@
Who is the child of Caro-
line LeRoy’s spouse?
Document 1::